OpenStudy (anonymous):

Given the function f(x) = 4 - 2x: If the domain is {-4, 0, 5}, find the range. A) {-4, 4, -6} B) {-4, 4, 14} C) {12, 4, -6} D) {12, 4, 14)

OpenStudy (texaschic101):

domain is all the x values while the range is all the y values. f(x) is also the same as y. f(x) = 4 - 2x.......if the domain (x)=-4 sub in -4 for all the x's f(-4)=4 -2(-4) f(-4)= 4 + 8 f(-4) = 12 so we know that if the domain (x) is-4 then the range 12 now to find out what the range is when the domain is 0, just sub in 0 for x. Continue doing this and you will have found your answer.